OpenStudy (anonymous):

Plz can anybody suggest me a book which will make math base good?. My base in math is very poor.

hero (hero):

Get the Handbook of Mathematics for Engineers

hero (hero):

Well, I guess it depends on what level you're on, but it's a great math reference book

OpenStudy (anonymous):

in 11

hero (hero):

These are all good questions...but if you're interested in ONE book, then the book I suggested is the one. If you have any doubts, go to a book store and peruse the book before buying it

OpenStudy (anonymous):

Thanks Hero can u suggest any book which will improve trignometry?

hero (hero):

I have a trig book. I'll just give it to you Whether it is good or not is relative to the individual. What one person might say is great, another person might not agree. But truthfully, whether or not any particular book resonates with you depends on your level.

OpenStudy (anonymous):

u r rit Hero

OpenStudy (anonymous):

r u sending the book?

hero (hero):

What you probably need is a good book PLUS someone to answer your questions as they arise. It will take me some time to create a link to the book. Once I have done so, I'll message it to you.

OpenStudy (anonymous):

Thanks Hero :)

OpenStudy (amistre64):

interactmath.com is a free resource with tons of practice problems from lots of textbooks. They even have the option to walk thru the solution process.

ganeshie8 (ganeshie8):

7 months have passed... how do u feel now ha :)

OpenStudy (anonymous):

need more practise

OpenStudy (ipwnbunnies):

Lol

ganeshie8 (ganeshie8):

my oly recommendation is, if you're spending time on learning stuff, spend it wisely, spend it on learning stuff correctly.. below are excellent resourses for a 12th grader : singel variable calculus : http://ocw.mit.edu/courses/mathematics/18-01-single-variable-calculus-fall-2006/video-lectures/ multivariable calculus : http://ocw.mit.edu/courses/mathematics/18-02-multivariable-calculus-fall-2007/video-lectures/ differential equations : http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/

ganeshie8 (ganeshie8):

it wud take atleast another year to do those courses and digest them... good luck ! :)

OpenStudy (fibonaccichick666):

I thoroughly recommend going to khan academy, they cover 1st grade to advanced in a relatively easy to understand form

OpenStudy (fibonaccichick666):

over a book that is

ganeshie8 (ganeshie8):

i like khanacademy for kindergarten to 12th grade levels i dont recommend it for higher levels

OpenStudy (fibonaccichick666):

it's good for calc

ganeshie8 (ganeshie8):

just personal opinion lol

ganeshie8 (ganeshie8):

MIT is a better place for calc

OpenStudy (fibonaccichick666):

lol I liked it up to calcs and some linear algebra

OpenStudy (fibonaccichick666):

but it's also a lot harder to understand

OpenStudy (fibonaccichick666):

khan is like the easy to understand version of mit lol

OpenStudy (anonymous):

lol

OpenStudy (fibonaccichick666):

but base i interpreted as basic, so that's why khan>mit ;)

ganeshie8 (ganeshie8):

i dont think so, MIT is the best in industry lol, khan is for range-of-the-moment pleasure folks... MIT is for long-term math pursuing ppl ;)

OpenStudy (fibonaccichick666):

ie basic vs detailed

OpenStudy (fibonaccichick666):

lol

OpenStudy (fibonaccichick666):

I thought this was for basic math?

ganeshie8 (ganeshie8):

khan is good for basic math < 12 ... non debatable :)

OpenStudy (accessdenied):

i guess my question would be, what does @Alivewidmusic need practice with at this time? any specifics?

OpenStudy (anonymous):

trig annoys me most of the time

OpenStudy (fibonaccichick666):

yea and mit is good for advanced topics non debatable :P

OpenStudy (accessdenied):

do you have an example problem that you've had trouble with? or maybe what topic is being covered that is causing you troubles.

OpenStudy (fibonaccichick666):

ok then yea, @Alivewidmusic you should be able to understand khanacademy.com They have some things in other languages too, but you can take their preliminary test, and they will place you so that you can practice.

OpenStudy (fibonaccichick666):

whoops it's khanacademy.org

OpenStudy (anonymous):

dont remember the prob.. bust its jus substituting problem answers get wrong

OpenStudy (accessdenied):

hm... something like this: \( \cos \left( \sin^{-1} \dfrac{a}{b} \right) \)? or using trig identities well? i don't think we're referring to the trig substitution for calc, which was my first idea of trig sub.

OpenStudy (anonymous):

yes similar to that

OpenStudy (accessdenied):

for problems much like that, my best tip is learning how to fill out the right triangles. really! a lot of those problems start by looking at the interior: \( \theta = \sin^{-1} \dfrac{a}{b} \) or \(\sin \theta = \dfrac{a}{b} \) by taking inverse of both sides. that last bit tells us the \(\text{opposite}\) length to \(\theta\) divided by the \(\text{hypotenuse}\) length. \( \sin \theta = \dfrac{a}{b} = \dfrac{\text{opposite}}{\text{hypotenuse}} \) so we can fill out two of three sides on a right triangle. the last one becomes available throguh pythagorean thm  and then we just have to find \(\cos \theta \) with the newly identified adjacent length. the same idea often permeates these sorts of problems. being able to draw and fill out all side lengths / possibly angles on the triangle should be powerful.

OpenStudy (anonymous):

hmm have to practise

OpenStudy (accessdenied):

i've been looking around a bit. it's hard to find this particular one (i didn't know what it would be called, so it took 4 searches). this looks like a good source of both guidance and practice problems: http://www.ck12.org/book/CK-12-Trigonometry-Concepts/r2/section/4.6/